Advanced Network Technology
University of Oregon
20070405 BGPd on eqix.routeviews.org crashed for unknown reasons. Down from about 11:50 to about 17:30 UTC. 20061211 After considering our options, we've decided to decommission route-view3 and manage all the Oregon and multi-hop sessions on either route-views, route-views2, or route-views6. All non-redundant sessions have been moved to one of the other collectors. 20061205 route-views3 is back up (with temporary hardware) and appears to be stable. 20061203 route-views3 is down with a hardware malfunction, and will be off line for repairs for an unknown period of time. If you are currently peering with route-views3 and are receiving alarms, please feel free to admin the sessions down until the problem has been resolved. 20061022 route-views2 was down from approximately 16:26 22-Oct-06 until 20:13 23-Oct-06 UTC due to an apparent hardware problem. 20060614 route-views.isc had an interface problem with caused collections to be lost from approximately 15:00 to 20:00 UTC. 20060505 route-views.eqix had problems of unknown origin for several hours. Some or all peer sessions were disrupted for part of this time. 20060413 The PAIX collector had interface problems, causing numerous peer resets throughout the day. A new interface was swapped in and, after a brief outage, connectivity appears to be restored. Problems were finally resolved prior to 23:00 UTC. 20060307 a fiber cut on campus caused outages of all the Oregon collectors (route-views, route-views2, and route-views3) between 03:30 and 05:00 UTC. 20060227 route-views2 down again briefly around 15:00 UTC. Between 16:30 and 16:45 the Zebra BGPD was replaced with a patched version. 20060226 route-views2 developed a problem with the Zebra process and was down from around 17:20 to 04:44 UTC. 20060213 route-views was down from approximately 20:30 to 21:40, then had to be reloaded again at 00:22 UTC. 20060213 routeviews.isc was down from approximately 15:30 to 00:30 UTC. 20060213 Ongoing issues with the BGP process on route-views.linx caused periodic loss of collections from approximately 27 Jan 06 to 10 Feb 06. We currently believe that the problem has now been fixed. 20060201 route-views.eqix was replaced with a new machine today. Resumption of service took longer than expected, and a large part of the day passed without any data being collected. 20051217 bgpd on route-views.linx died and had to be manually restarted, resulting in a loss of data from approximately 13:12 to 15:20 UTC. 20051122 memory problems on route-views caused a loss of connectivity and and data collection problems between 07:00 and 16:00 UTC. 20050809 a switch failure caused loss of connectivity to the majority of peers from approximately 22:15 to 22:45 UTC today. This will show up as a loss of data in the archive files for today. 20050405 route-views2 ran out of disk space. Approximately 2 hours of collection was lost. 20050328 through 20050330 route-views.eqix continues to experience hardware problems, and is down for about two days. We are still waiting for new hardware. 20050308 route-views.oregon-ix stopped accepting tcp connections and had to be rebooted. Cause of the problem is unknown. 20050201 route-views.eqix was down for approximately a day and a half. This machine is experiencing ongoing problems and a replacement is being ordered. 20041021 upstream router crash affected several BGP peers on the boxes at the University. 20041012 route-views.eqix, hung and had to be rebooted. Cause unknown. 20041005 route-views.isc, NAP switch outage caused a few hours of peer losage. 20040927 route-views.eqix ethernet PHY wedged. rebooted 20040911 route-views.isc (PAIX) NAP connection affected by UPS failure. 20040817 route-views3 routing daemon restarted. It became very confused after adding a specific hold-timer to the BGP groups. 20040810 route-views.isc (PAIX) NAP interface accumulating errors that are affecting BGP sessions. Began about 12:45 UTC. Thought to be an interface problem, box was rebooted. Later found that S&D is having switch problems. 20040715 route-views2 rebooted at 19:00. Suspected new kernel as cause of lingering tcp connections. Not the case, but did clear the result of these in zebra. 20040702 route-views.linx rebooted at 01:00 to load new kernel. 20040413 route-views2 dropping sessions during RIB dumps after upgrade. Addition of memory, fixing RAID microcode, and code changes seem to have fixed the problem. 20040623 route-views.wide lost power at ~10:25 UTC 20040405 route-views2 and 6 upgraded. 20040329 route-views.isc lost power. 20040324 route-views.linx rebooted to pick-up new software. route-views.wide rebooted to pick-up new software. 20040303 route-views.wide now has a connection to the WIDE/NSPIXP2 IPv6 switch and is ready for IPv6 peering. 20040225 route-views.wide.routeviews.org lost power early this morning (PDT). Cause not yet known. 20040223 We had additional problems with the archive.routeviews.org hardware. Nearly all is restored now. 20031212 We suffered a RAID failure on archive.routeviews.org, the machine that also collects RIBs from the Cisco route-views.routeviews.org. It successfully rebuilt on the 12th and failed again the morning of the 13th. It has been rehomed to a new machine with a more reliable RAID controller. We regret that Cisco RIBs that would have been collected during these periods are lost.